OpenStudy (anonymous):

Dave has 7 coins. He only has pennies,nickels, and dimes. He has more nickels then pennies and more dime then nickels. How much money does he have?

OpenStudy (jamesj):

So let P be the number of pennies, N = number of nickels, D = # of dimes. Then P + N + D = 7 and D > N > P what values can P, N and D have?

OpenStudy (anonymous):

This is a thrid grade math sheet so i typed the question as it was worded on the sheet.

OpenStudy (turingtest):

x=#pennies y=#nickels z=#dimes x+y+z=7 z>y>x z=4 y=2 x=1 works that's 4(0.1)+2(0.05)+1(0.01)=$0.51